### v1

@@ V-C Generalization
V-D Ablation Study
Different methods for generating poisoned texts. To ensure the effectiveness of the evaluation, we employ three different methods introduced by PoisonedRAG, GARAG, and PRCAP to generate the poisoned texts. The experimental results in Table V show that RevPRAG consistently achieves high TPRs and low FPRs when confronted with poisoned texts generated by different strategies. For instance, RevPRAG achieved 97.2% (with GPT2-XL 1.5B), 98.5% (with Llama2-7B), and 97.7% (with Mistral-7B) TPRs for poisoned texts generated with PoisonedRAG. Our method focuses on detecting the outputs of RAG rather than analyzing the inputs. In other words, regardless of the content of inputs into LLM or the poisoned text generation strategy, our method can reliably detect poisoned responses as long as they occur.
@@ V-D Ablation Study
Quantity of injected poisoned texts. Table VI illustrates the impact of varying quantities of poisoned text on the detection performance of RevPRAG. Injecting different amounts of poisoned text into the knowledge database impacts the likelihood of retrieving poisoned content. The more poisoned texts are injected, the higher the likelihood of retrieving them for RAG processing. This is because one of the optimization objectives for generating poisoned text is to maximize its similarity to the target query. In our experiments, we find that when five poisoned texts are injected into the knowledge database, there is a high likelihood of retrieving all five poisoned texts when the total number of retrieved texts is also five (following practice of prior work). From the experimental results, we observe that even with varying amounts of injected poisoned text, RevPRAG consistently achieves high TPRs and low FPRs. For example, when the total number of retrieved texts is five and the injected quantity is two, RevPRAG achieves a 99.5% TPR and a 4.7% FPR for RAG with Llama2-7B. The reason for this phenomenon is that the similarity between the retrieved poisoned texts and the query is higher than that of clean texts. Consequently, the LLM generates responses based on the content of the poisoned texts.
TABLE VII: Impact of similarity metric.
@@
Impact of similarity metric. Different methods for calculating similarity between embedding vectors of queries and texts in the knowledge database may lead to varying poisoning effects and distinct LLM activations. Therefore, it is crucial to conduct ablation experiments using various similarity metrics. Table VII shows the results on the HotpotQA dataset, indicating that the choice of similarity calculation method has minimal impact on RevPRAG’s performance, which consistently achieves high TPR and low FPR. This suggests the robustness of our approach, as it reliably identifies poisoned texts even when LLM activations vary slightly under similar conditions.
(a) TPRs of RevPRAG.
(b) FPRs of RevPRAG.
Figure 6: Activations from specified layers.
Activations from specified layers. Fig. 6 illustrates the detection performance of RevPRAG using activations from different layers of various LLMs on HotpotQA. In our approach, we utilize activations from all layers as both training and testing data, yielding excellent results. Additionally, the experimental results demonstrate that utilizing activations from only the first few layers can still achieve satisfactory detection performance, providing valuable insights for future research. For example, when using activations from layers 0 to 5, RevPRAG achieved TPRs exceeding 97% while maintaining FPRs below 7% for RAG with all LLMs on HotpotQA. Furthermore, the experimental results suggest that using activations from intermediate or deeper layers can lead to performance fluctuations, including signs of degradation or slower convergence. For instance, when using activations from layers 16 to 24 with Llama3-8B as the LLM in RAG, RevPRAG achieves a TPR of 78.8% on NQ dataset and 86% on MS-MARCO dataset.
Effects of different support set size. We experiment with various support set sizes ranging from 50 to 250 to examine their effect on the performance of RevPRAG. This evaluation was conducted on Llama2-7B with different datasets. The results in Fig. 7 indicate that varying the support size does not significantly impact the model’s detection performance.
Figure 7: Effects of support set size.
V-E Isolating Poisoned Responses and Hallucinations
It is well-known that hallucinations are an inevitable phenomenon in LLMs. Even with the introduction of a knowledge database in RAG, LLMs may still generate non-factual responses due to hallucinations. Therefore, the incorrect responses generated by RAG may also stem from hallucinations, rather than being solely caused by RAG poisoning. Fig. 8 shows the t-SNE [58] representation of mean activations for poisoned response and hallucinations across all layers for Mistral-7B and Llama2-7B on the NQ dataset. We observe that activations across all layers clearly distinguish between hallucinations and poisoned responses. This key finding has led us to extend our approach to differentiate between poisoned responses and hallucinations.
@@ V-E Isolating Poisoned Responses and Hallucinations
TABLE VIII: RevPRAG could achieve high TPRs and low FPRs for distinguishing between poisoned responses and hallucinations.
@@
We continue to collect data and train the model using the process outlined in Fig. 2, with the only difference being that we now collect hallucination data. We also conduct extensive experiments on RAG with different LLMs and datasets. From the experimental results in Table VIII, we can see that our method achieves a high True Positive Rate (TPR) across all LLMs and datasets. For instance, RevPRAG achieved 98.7% (on NQ), 97.5% (on HotpotQA), and 97.3% (on MS-MARCO) TPRs for RAG with GPT2-XL 1.5B. Furthermore, we observe that the FPR remains low across all evaluation settings. As shown in the table, RevPRAG could achieve 0.8% (on NQ), 0.8% (on HotpotQA) and 0.6% (on MS-MARCO) FPRs for RAG with Llama3-8B. This further supports our previous observation that there is a clear distinction between poisoned responses and hallucinations.
To the best of our knowledge, this is the first work that successfully achieves to effectively differentiate poisoned response from hallucinations.
VI Conclusion
In this work, we find that correct and poisoned responses in RAG exhibit distinct differences in LLM activations. Building on this insight, we develop RevPRAG, a detection pipeline that leverages these activation patterns to identify poisoned responses in RAG caused by the injection of malicious text into the knowledge database. Our approach demonstrates robust performance across RAGs utilizing five different LLMs and four distinct retrievers on three datasets. Experimental results show that RevPRAG achieves exceptional accuracy, with true positive rates approaching 98% and false positive rates near 1%. Ablation studies further confirm its effectiveness in detecting poisoned responses under various poisoning attack types and intensities. Additionally, we validate our method’s ability to distinguish poisoned responses from non-factual responses induced by hallucinations, ensuring precise identification of genuinely poisoned outputs. Overall, our approach reliably detects poisoned responses without disrupting the functionality of RAG systems.
VII Acknowledgments
The authors affiliated with Fudan University are supported by the National Key R&D Program of China 2023YFB3107404.

### v2

@@ 5.3 Ablation Study
Quantity of injected poisoned texts. Table 5 illustrates the impact of varying quantities of poisoned text on the detection performance of RevPRAG. The more poisoned texts are injected, the higher the likelihood of retrieving them for RAG processing. From the experimental results, we observe that even with varying amounts of injected poisoned text, RevPRAG consistently achieves high TPRs and low FPRs. For example, when the total number of retrieved texts is five and the injected quantity is two, RevPRAG achieves a 99.5% TPR and a 4.7% FPR for RAG with Llama2-7B. The reason for this phenomenon is that the similarity between the retrieved poisoned texts and the query is higher than that of clean texts. Consequently, the LLM generates responses based on the content of the poisoned texts.
Effects of different support set size. In RevPRAG, the Support Data plays a pivotal role in the model’s learning process. It supplies labeled information and task-specific knowledge, allowing the model to conduct effective reasoning and learning even with limited data. We experiment with various support set sizes ranging from 50 to 250 to examine their effect on the performance of RevPRAG. This evaluation was conducted on Llama2-7B with different datasets. The results in Fig. 6 indicate that varying the support size does not significantly impact the model’s detection performance.
Figure 6: Effects of support set size.
6 Conclusion
In this work, we find that correct and poisoned responses in RAG exhibit distinct differences in LLM activations. Building on this insight, we develop RevPRAG, a detection pipeline that leverages these activations to identify poisoned responses in RAG caused by the injection of malicious texts into the knowledge database. Our approach demonstrates robust performance across RAGs utilizing five different LLMs and four distinct retrievers on three datasets. Experimental results show that RevPRAG achieves exceptional accuracy, with true positive rates approaching 98% and false positive rates near 1%. Ablation studies further validate its effectiveness in detecting poisoned responses across different types and levels of poisoning attacks. Overall, our approach can accurately distinguish between correct and poisoned responses.
Limitations.
Our work has the following limitations:
  • This work does not propose a specific method for defending against poisoning attacks on RAG. Instead, our focus is on the timely detection of poisoned responses generated by the LLM, aiming to prevent potential harm to users from such attacks.
  • Our approach requires accessing the activations of the LLM, which necessitates the LLM being a white-box model. While this may present certain limitations for users, our method can be widely adopted by LLM service providers. Providers can implement our strategy to ensure the reliability of their services and enhance trust with their users.
  • Our approach primarily focuses on determining whether the response generated by the RAG is correct or poisoned, without delving into more granular distinctions. The main goal of our study is to protect users from the impact of RAG poisoning attacks, while more detailed classifications of RAG responses will be addressed in future work.
Ethics Statement
The goal of this work is to detect whether a RAG has generated a poisoned response. All the data used in this study is publicly available, so it does not introduce additional privacy concerns. All source code and software will be made open-source. While the open-source nature of the code may lead to adaptive attacks, we can further enhance our model by incorporating more internal and external information. Overall, we believe our approach can further promote the secure application of RAG.
Appendix A Additional Results
Table 6: Generalization Performance of RevPRAG for RAG with four different LLMs. The training and test datasets vary across different rows. Abbreviations: Hot (HotpotQA), MS (MS-MARCO).
Training Dataset Test Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
NQ & Hot MS TPR 0.881 0.886 0.948 0.956
FPR 0.134 0.149 0.076 0.066
Hot & MS NQ TPR 0.980 0.983 0.988 0.980
FPR 0.007 0.074 0.078 0.038
NQ & MS Hot TPR 0.977 0.961 0.942 0.978
FPR 0.025 0.089 0.055 0.049
NQ & Hot & MS NQ & Hot & MS TPR 0.986 0.994 0.985 0.987
FPR 0.032 0.007 0.009 0.035
A.1 Generalization
Given the wide range of RAG application scenarios and the diverse user requirements it faces, it is impractical to ensure that our detection model has been trained on all possible scenarios and queries in real-world applications. However, the performance of neural network models largely depends on the similarity between the distributions of the training data and the test data Yang et al. (2024). Consequently, our model’s performance may degrade when faced with training and test data that stem from differing distributions—a challenge frequently observed in real-world scenarios. To address this issue, we conduct a series of generalization experiments. Specifically, we train the detection model using any two datasets and test it on a third dataset that was not used during training. For example, we use NQ and HotpotQA as training datasets and MS-MARCO as the testing dataset. Although these three datasets are all QA datasets, they exhibit significant differences. For example, NQ focuses on extracting answers to factual questions from a single long document, HotpotQA involves multi-document reasoning to derive answers, and MS-MARCO retrieves and ranks relevant answers from a large-scale collection of documents. Therefore, conducting generalization experiments based on these three datasets is reasonable.
Table 6 illustrates the TPRs and FPRs of RevPRAG for RAG with four different LLMs. Overall, the experimental results demonstrate that our detection model exhibits strong generalization performance across RAG with different LLMs and various datasets. For example, when using HotpotQA and MS-MARCO as training data, the detection model achieves TPRs of 98% (with GPT2-XL 1.5B), 98.3% (with Llama2-7B), 98.8% (with Mistral-7B), and 98% (with Llama3-8B) on the NQ dataset. Meanwhile, all FPRs remain below 8%. Furthermore, we observe that the generalization performance is best when NQ is used as the test data (for instance, 98.3% with Llama2-7B), while MS-MARCO shows the poorest performance (for instance, 88.6% with Llama2-7B). We attribute this to the fact that the questions and tasks in HotpotQA and MS-MARCO are more complex compared to those in NQ. Therefore, detection models trained on more complex tasks generalize well to simpler tasks, whereas the reverse is more challenging. In conclusion, these experimental results highlight that RevPRAG exhibits strong generalization and robust detection performance, even in the presence of significant discrepancies between the training and test datasets.
(a) TPRs of RevPRAG.
(b) FPRs of RevPRAG.
Figure 7: RevPRAG trained on the activations from specific layers.
A.2 RevPRAG’s Performance on Complex Open-Ended Questions
In this section, we conducted a series of experiments to evaluate the performance of RevPRAG on complex, open-ended questions (e.g., “how to make relationship last?”). These questions present unique challenges due to their diverse and unstructured nature, in contrast to straightforward, closed-ended questions (e.g., “What is the name of the highest mountain?”). In our experiments, the NQ, HotpotQA, and MS-MARCO datasets primarily consist of close-ended questions. As a result, the majority of our previous experiments focused on close-ended problems, which was our default experimental setting. In this study, we utilized the advanced GPT-4o to filter and extract 3,000 open-ended questions from the HotpotQA and MS-MARCO datasets for training and testing the model. For open-ended questions, cosine similarity is employed to evaluate whether the LLM’s response aligns with the attacker’s target response. If the similarity surpasses a predefined threshold, it is considered indicative of a successful poisoning attack.
The experimental results are shown in Table 7. We can observe that RevPRAG demonstrates excellent detection performance even on complex open-ended questions. For example, RevPRAG achieved TPRs of 99.1% on HotpotQA and 99.0% on MS-MARCO, alongside FPRs of 0.8% on HotpotQA and 0.1% on MS-MARCO for RAG utilizing the Mistral-7B model.
Table 7: RevPRAG achieved high TPRs and low FPRs on the open-ended questions from HotpotQA and MS-MARCO datasets.
Dataset Metrics LLMs of RAG
GPT2-XL 1.5B Llama2-7B Mistral-7B Llama3-8B
HotpotQA TPR 0.982 0.995 0.991 0.982
FPR 0.033 0.029 0.008 0.007
MS-MARCO TPR 0.988 0.989 0.990 0.983
FPR 0.009 0.009 0.001 0.017
Table 8: RevPRAG trained on the activations from specific individual layers of GPT2-XL 1.5B.
Dataset Metrics Different layers
layer 0 layer 8 layer 15 layer 22 layer 29 layer 36 layer 41 layer 47
NQ TPR 0.996 0.988 0.996 0.984 0.996 0.988 0.992 0.996
FPR 0.027 0.007 0.017 0.003 0.007 0.003 0.017 0.003
HotpotQA TPR 0.713 0.984 0.994 0.989 0.520 0.619 0.931 0.992
FPR 0.409 0.023 0.012 0.006 0.445 0.409 0.023 0.019
MS-MARCO TPR 0.967 0.998 0.988 0.986 0.988 0.963 0.955 0.992
FPR 0.023 0.004 0.002 0.019 0.030 0.026 0.037 0.017
A.3 Activations from Specified Layers.
Fig. 7 illustrates the detection performance of RevPRAG using activations from different layers of various LLMs. In previously presented experiments, we utilize activations from all layers as both training and testing data, yielding excellent results. Additionally, we also test using different layers. The experimental results in Fig. 7 demonstrate that utilizing activations from only the first few layers can still achieve satisfactory detection performance, providing valuable insights for future research. For example, when using activations from layers 0 to 5, RevPRAG achieved TPRs exceeding 97% while maintaining FPRs below 7% for RAG with all LLMs on HotpotQA. However, the experimental results also suggest that using activations from intermediate or deeper layers can lead to performance fluctuations, including signs of degradation or slower convergence. For instance, when using activations from layers 16 to 24 with Llama3-8B as the LLM in RAG, RevPRAG achieves a TPR of 78.8% on NQ dataset and 86% on MS-MARCO dataset.
We further explored the use of activations from a specific individual layer of the LLMs to train and test RevPRAG. We chose 8 layers with roughly even spacing for testing. As shown in Table 8, when using activations from only a specific layer of the GPT2-XL model, RevPRAG demonstrates excellent performance in general. For instance, when the model is trained using activations from layer 0 on the NQ dataset, the TPR can reach as high as 99.6%. However, we also observed that activations from certain layers do not yield satisfactory performance. For example, when the model is trained using activations from layer 29 on the HotpotQA dataset, the TPR is only 52%, while the FPR reaches 44.5%. It is precisely due to the existence of these suboptimal layers that models trained with multi-layer activations may not always outperform those using single-layer activations (such as layer 0 with NQ dataset). However, incorporating multi-layer activations can enhance the model’s stability, mitigating the detrimental effects of these suboptimal layers.
A.4 Impact of Similarity Metric.
Table 9 shows the results on the HotpotQA dataset, indicating that the choice of similarity calculation method has minimal impact on RevPRAG’s performance, which consistently achieves high TPR and low FPR. This suggests the robustness of our approach, as it reliably identifies poisoned texts even when LLM activations vary slightly under similar conditions.
Table 9: Impact of similarity metric.
@@ A.2 RevPRAG’s Performance on Complex Open-Ended Questions
A.5 Isolating Poisoned Responses and Hallucinations
It is well-known that hallucinations are an inevitable phenomenon in LLMs. Even with the introduction of a knowledge database in RAG, LLMs may still generate non-factual responses due to hallucinations. Therefore, the incorrect responses generated by RAG may also stem from hallucinations, rather than being solely caused by RAG poisoning. We conducted experiments to test if our approach can distinguish hallucinations and RAG poisoning. Fig. 8 shows the t-SNE representation of mean activations for poisoned response and hallucinations across all layers for Mistral-7B and Llama2-7B on the NQ dataset. We observe that activations across all layers clearly distinguish between hallucinations and poisoned responses.
@@ A.5 Isolating Poisoned Responses and Hallucinations
This key finding has led us to extend our approach to differentiate between poisoned responses and hallucinations. We thus continue to collect data and train the model using the process outlined in Fig. 3, with the only difference being that we now collect hallucination data. We also conduct extensive experiments on RAG with different LLMs and datasets. From the experimental results in Table 10, we can see that our method achieves a high TPR across all LLMs and datasets. For instance, RevPRAG achieved 98.7% (on NQ), 97.5% (on HotpotQA), and 97.3% (on MS-MARCO) TPRs for RAG with GPT2-XL 1.5B. Furthermore, we observe that the FPR remains low across all evaluation settings. As shown in the table, RevPRAG could achieve 0.8% (on NQ), 0.8% (on HotpotQA) and 0.6% (on MS-MARCO) FPRs for RAG with Llama3-8B. This further supports our previous observation that there is a clear distinction between poisoned responses and hallucinations.
Table 10: RevPRAG could achieve high TPRs and low FPRs to distinguish poisoned responses and hallucinations.